An investor in Treasury securities expects inflation to be 2.5% in Year 1, 3.1% in Year 2, and 4.4% each year thereafter. Assume that the real risk-free rate is 1.55% and that this rate will remain constant. Three-year Treasury securities yield 6.25%, while 5-year Treasury securities yield 7.95%. What is the difference in the maturity risk premiums (MRPs) on the two securities; that is, what is MRP5 - MRP3? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Step 1 - Inflation risk premium for 3 year and 5 year:

Inflation premium (3 year) = (2.5 + 3.1 + 4.4)/3 = 3.33%

Inflation premium (5 year) = (2.5 + 3.1 + 4.4 + 4.4 +4.4)/5 = 3.76%

The real risk free rate is 1.55%

MRP3 = 6.25% - 3.33% - 1.55% = 1.37%

MRP5 = 7.95% - 3.76% -1.55% = 2.64%.

So, MRP5 - MRP3 = 1.27%